OpenStudy (anonymous):

PLEASE HELP!! MEDALS!! The function f(x) = 16(3)x represents the growth of a dragonfly population every year in a remote swamp. Erin wants to manipulate the formula to an equivalent form that calculates four times a year, not just once a year. Which function is correct for Erin's purpose, and what is the new growth rate?

OpenStudy (anonymous):

f(x) = 4(3)x; growth rate 300% f(x) = 4(3)x, growth rate 4% f(x) = 4(1.32)x; growth rate 4% f(x) = 4(1.32)4x; growth rate 132%

OpenStudy (anonymous):

everything is ^x not multiplied

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@masterkman

OpenStudy (anonymous):

Spamming doesn't help you.

OpenStudy (anonymous):

Spamming?

OpenStudy (anonymous):

I'm justmentioning people that could help me

OpenStudy (anonymous):

Wait patiently like everyone else.

OpenStudy (campbell_st):

is the growth formula \[f(x) = 16 \times 3^x\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks for the answer btw!

OpenStudy (campbell_st):

ok... well the problem is that the formula doesn't make a lot of sense... is the growth rate 16% per year...?

OpenStudy (anonymous):

Yes. I think haha we can go with that

OpenStudy (campbell_st):

ok.... well I'd say you divide the growth rate by 4 and then increase the number of time periods my multiplying by 4 is there any other information in the question...?

OpenStudy (anonymous):

Nope! thats it! I hate these online courses!

OpenStudy (campbell_st):

wow its a badly written question... my best guess is the last option...

OpenStudy (anonymous):

I got it right! thanks so much!

OpenStudy (campbell_st):

glad to help